# DocMath-Eval: Evaluating Math Reasoning Capabilities of LLMs in Understanding Long and Specialized Documents

### v2

@@ Appendix A Appendix
1 | 1 working and 1 internship at US | New subset, Annotation validation |
2 | >= 2 internship at US | New subset, Annotation validation |
3 | 1 working at Singapore and 2 internship at US | New subset |
4 | 2 working and >= 1 internship at US | New subset |
5 | 1 internship at US, 2 internship at China | Re-annotation on three subsets, Annotation validation |
6 | Graduate student majored in computer science | Re-annotation on three subsets, Annotation validation |
7 | Graduate student majored in statistics | Re-annotation on three subsets |
Model | Organization | Size | Notes | Source |
GPT-4-Turbo | OpenAI | – | gpt-4o-2024-05-13 | |
GPT-4o | OpenAI | – | gpt-4-turbo-2024-04-09 | |
GPT-3.5-Turbo | OpenAI | – | gpt-3.5-turbo-0125 | |
@@
Claude-3-Opus | Anthropic | – | claude-3-opus-20240229 | |
Claude-3-Sonnet | Anthropic | – | claude-3-sonnet-20240229 | |
Claude-3-Haiku | Anthropic | – | claude-3-haiku-20240307 | |
@@
Qwen2 | Alibaba | 7 & 72B | Qwen/Qwen2-*B-Instruct | |
@@
Llama-3 | Meta | 8 & 70B | meta-llama/Meta-Llama-3-*B-Instruct | |
Llama-3.1 | Meta | 8 & 70B & 405B | meta-llama/Meta-Llama-3.1-*B-Instruct | |
@@
Mistral-Nemo | Mistral AI | 12B | mistralai/Mistral-Nemo-Instruct-2407 | |
Mistral-Large | Mistral AI | 123B | mistralai/Mistral-Large-Instruct-2407 | |
Mathstral | Mistral AI | 7B | Math-Specific | mistralai/Mathstral-7B-v0.1 |
Mixtral | Mistral AI | 46 & 141B | MoE | mistralai/Mixtral--Instruct-v0.1 |
Codestral | Mistral AI | 22B | Code-Specific | mistralai/Codestral-22B-v0.1 |
@@
DeepSeek-Coder-V1 | DeepSeek | 33B | Code-Specific | deepseek-ai/deepseek-coder-33b-instruct |
DeepSeek-V2 | DeepSeek | 16 & 236B | MoE | deepseek-ai/DeepSeek-V2-Lite-Chat. We use the official API provided by DeepSeek for deepseek-ai/DeepSeek-V2-Chat |
DeepSeek-Coder-V2 | DeepSeek | 16 & 236B | MoE | deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct. We use the official API provided by DeepSeek for deepseek-ai/DeepSeek-Coder-V2-Instruct |
@@
Phi-3-Mini | Microsoft | 3B | microsoft/Phi-3-mini-4k-instruct | |
@@
InternLM2-Math-Plus | InternLM | 7B | Math-Specific | internlm/internlm2-math-plus-7b |
@@
WizardMath | WizardLM Team | 7B | Math-Specific | WizardLMTeam/WizardMath-7B-V1.1 |
WizardCoder | WizardLM Team | 33B | Code-Specific | WizardLMTeam/WizardCoder-33B-V1.1 |
WizardLM-2 (MoE) | WizardLM Team | 141B | MoE | alpindale/WizardLM-2-8x22B |
[…]
@@
DeepSeek-Coder-V2 | 236B | Code | 88.2 | 80.9 | 74.9 | 67.2 | 56.8 | 54.8 | 37.5 | 31.5 | 61.9 | 55.7 | ||||

### v3

@@ Appendix A Appendix
| 1 | 1 working and 1 internship at US | New subset, Annotation validation |
| 2 | >= 2 internship at US | New subset, Annotation validation |
| 3 | 1 working at Singapore and 2 internship at US | New subset |
| 4 | 2 working and >= 1 internship at US | New subset |
| 5 | 1 internship at US, 2 internship at China | Re-annotation on three subsets, Annotation validation |
| 6 | Graduate student majored in computer science | Re-annotation on three subsets, Annotation validation |
| 7 | Graduate student majored in statistics | Re-annotation on three subsets |
| Organization | Model | Size | Notes | Source |
| OpenAI | GPT-4-Turbo | – | gpt-4o-2024-05-13 | |
| GPT-4o | – | gpt-4-turbo-2024-04-09 | ||
| GPT-3.5-Turbo | – | gpt-3.5-turbo-0125 | ||
@@
| Claude-3-Opus | – | claude-3-opus-20240229 | ||
| Claude-3-Sonnet | – | claude-3-sonnet-20240229 | ||
| Claude-3-Haiku | – | claude-3-haiku-20240307 | ||
@@
| Alibaba | Qwen2 | 7 & 72B | Qwen/Qwen2-*B-Instruct | |
@@
| Llama-3 | 8 & 70B | meta-llama/Meta-Llama-3-*B-Instruct | ||
| Llama-3.1 | 8 & 70B & 405B | meta-llama/Meta-Llama-3.1-*B-Instruct | ||
@@
| Mistral-Nemo | 12B | mistralai/Mistral-Nemo-Instruct-2407 | ||
| Mistral-Large | 123B | mistralai/Mistral-Large-Instruct-2407 | ||
| Mathstral | 7B | Math-Specific | mistralai/Mathstral-7B-v0.1 | |
| Mixtral | 46 & 141B | MoE | mistralai/Mixtral--Instruct-v0.1 | |
| Codestral | 22B | Code-Specific | mistralai/Codestral-22B-v0.1 | |
@@
| DeepSeek-Coder-V1 | 33B | Code-Specific | deepseek-ai/deepseek-coder-33b-instruct | |
| DeepSeek-V2 | 16 & 236B | MoE | deepseek-ai/DeepSeek-V2-*-Chat | |
| DeepSeek-Coder-V2 | 16 & 236B | Code-Specific, MoE | deepseek-ai/DeepSeek-Coder-V2-*-Instruct | |
@@
| Phi-3-Mini | 3B | microsoft/Phi-3-mini-4k-instruct | ||
@@
| InternLM2-Math-Plus | 7B | Math-Specific | internlm/internlm2-math-plus-7b | |
@@
| WizardMath | 7B | Math-Specific | WizardLMTeam/WizardMath-7B-V1.1 | |
| WizardCoder | 33B | Code-Specific | WizardLMTeam/WizardCoder-33B-V1.1 | |
| WizardLM-2 (MoE) | 141B | MoE | alpindale/WizardLM-2-8x22B | |
@@
| DeepSeek-Coder-V2 | 236B | Code, MoE | 88.2 | 80.9 | 74.9 | 67.2 | 56.8 | 54.8 | 37.5 | 31.5 | 61.9 | 55.7 | ||||